OpenStudy (anonymous):

Find the projection of u onto v. Then write u as the sum of two orthogonal vectors, one of which is projvu . 1. u = <3, –4> v = <2, 1>

OpenStudy (anonymous):

Do you remember the formula for projection?

OpenStudy (anonymous):

no

OpenStudy (anonymous):

thats why i don't know how to solve the problem

OpenStudy (anonymous):

The unit vector of the one being projected on, times the comp

OpenStudy (anonymous):

ok thanks :D

OpenStudy (anonymous):

Do you know the comp?

OpenStudy (anonymous):

lol i will figure it out

OpenStudy (anonymous):

It is the dot product divided by the magnitude of he vector being projected onto

OpenStudy (anonymous):

\[ \text{proj}_\mathbf{v}\mathbf{u}=\text{comp}_\mathbf{v}\mathbf{u}\frac{\mathbf{v}}{||\mathbf{v}||} = \frac{\mathbf{v}\cdot \mathbf{u}}{||\mathbf{v}||}\frac{\mathbf{v}}{||\mathbf{v}||} = \frac{\mathbf{v}\cdot \mathbf{u}}{||\mathbf{v}||^2}\mathbf{v} = \frac{\mathbf{v}\cdot \mathbf{u}}{\mathbf{v}\cdot \mathbf{v}}\mathbf{v} \]

OpenStudy (anonymous):

never mind. i was doing the wrong homework. no wonder i didn't understand it